After how many minutes from 12:00 o’clock noon will the

minute hand and the hour hand of the clock makes and

angle of 120degrees for the first time.


1
Expert's answer
2021-09-20T16:36:09-0400

Let x = position of the minute hand (in minutes) when this occurs.

The minute hand travels 360 degrees in 1 hr

The hr hand travels "\frac {360} {12}" = 30 degrees, in 1 hr

The fraction of an hr can be expressed by "\frac {x} {60}"

Min hand degrees - hr hand degrees = 120

"(\frac {x} {60})×360-(\frac {x} {360}) ×30=120"

"6x - 5x = 120"

"5.5x = 120"

"x = \frac {120} {5.5}"

"x" = 21.82 minutes or 21 min 49 sec